following are two weekly forecasts made by two different methods for the number of gallons of gasoline, in thousands, demanded at a local gasoline station. Also shown are actual demand levels in thousands of gallons:

Forecasts

week Method 1 Method 2 Actual Demand

1 0.90 0.80 0.70

2 1.05 1.20 1.00

3 0.95 0.90 1.00

4 1.20 1.11 1.00

The MAD method 1 = ??? thousand gallons... 2. The absolute deviatin based on the forecast developed using Method 1 adds to ??? thousand gallons. Mean squared error (MSE) is the average of (actual-Forecast). from the information given in method 1 , the value of n=4. the value E(autal-forecast) will be??? thousand gallons.
Answer #1

Method 1

Week

Method 1

Actual

E

E^2

1

0.9

0.7

0.2

0.04

2

1.05

1

0.05

0.0025

3

0.95

1

0.05

0.0025

4

1.2

1

0.2

0.04

0.5

0.085

MAD = sum of error/n

           = 0.5/4

           =0.125

MSE = sum of squared error/n

         = 0.0.85/4

         = 0.02125

Method 2

Week

Method 2

Actual

E

E^2

1

0.8

0.7

0.1

0.01

2

1.2

1

0.2

0.04

3

0.9

1

0.1

0.01

4

1.11

1

0.11

0.0121

0.51

0.0721

MAD = 0.51/4 = 0.1275

MSE = 0.0721/4 = 0.018